OpenStudy (anonymous):

What is the equation of the line that passes through y-(-3)=M(x-(-2) and is perpendicular to 2x–3y = 6? can someone please help me to understand this. not just giving me the answer but teaching me how to do it myself.

OpenStudy (loser66):

ready?

OpenStudy (loser66):

no? ok, hihihih,..

OpenStudy (anonymous):

yes please, Mr. Mumble lol

OpenStudy (loser66):

first off, your problem is unclear about the point. What does y-(-3)=M(x-(-2) mean?

OpenStudy (anonymous):

it is in slope-intercept form originally it is the point (-2,-3)

OpenStudy (loser66):

Thanks for that. Let it there. I start with the point M (-2,-3) and the given line, OK?

OpenStudy (loser66):

the given line is 2x -3y =6 , isolate y, what do you have?

OpenStudy (anonymous):

M just ='s the slope of the line which is unknown.

OpenStudy (loser66):

Forget it, just start with the point and the given line. Period. Work, and show me what you have after isolate y on the given line.

OpenStudy (anonymous):